OpenStudy (anonymous):

Proof: sin t (csc t - sin t)= cos ^ 2 t

OpenStudy (anonymous):

Could you show us what you've done so far please?

OpenStudy (anonymous):

No I don't know how to do it. We have to simplify the left side to match the right

OpenStudy (anonymous):

Yes, that's a start. Why don't you start off by expanding/distributing the LHS for me? You could get rid fo the brackets/paranthesis.

OpenStudy (anonymous):

rid of*

OpenStudy (anonymous):

Idk how to do that

OpenStudy (anonymous):

First, you have to pick the most complicated side, sin(csct-sint) right? sin t (csc t - sin t)= cos ^ 2 t that means you have to do.. sintcsct-sintsint=cos^2t because you're distributing now.. you can change csct because it equals 1/sint so.. (sint)(1/sint)-sintsint=cos^2t the sin cancel and you're left with 1. sintsint turns into sin^2t sin^2t equals.... 1-cos^2t so 1-1-cos^2t the ones cancel out so you're left with... cos^2t = cos^2t